# Manipulating Feature Visualizations with Gradient Slingshots

### v3

@@ D.1 Accuracy – Manipulation Trade-Off
We expand on the experimental results of the accuracy–manipulation trade-off analysis for the Gradient Slingshots attack, introduced in Sec. 4.1, with additional findings reported in Tables 13, 14, 15, 16, 10, 11, 12 and 13
@@
Table 13: Accuracy–manipulation trade-off for the Gradient Slingshots attack on the “zero” output neuron in 6-layer CNN . Reported are test accuracy (in %), AUROC for the “zero” class, and the mean ± standard deviation of similarity metrics between the manipulated FV and the target image.
$\alpha$ Acc. AUROC CLIP $\uparrow$ MSE $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$
Original 99.67 1.00 $0.88\pm 0.01$ $0.13\pm 0.01$ $0.15\pm 0.01$ $0.04\pm 0.03$
0.900000 99.28 1.00 $0.89\pm 0.01$ $0.10\pm 0.00$ $0.15\pm 0.01$ $0.13\pm 0.02$
0.800000 98.98 1.00 $0.93\pm 0.01$ $0.03\pm 0.00$ $0.11\pm 0.02$ $0.74\pm 0.01$
0.700000 98.89 1.00 $0.95\pm 0.01$ $0.02\pm 0.00$ $0.07\pm 0.02$ $0.76\pm 0.03$
0.500000 98.10 1.00 $0.86\pm 0.01$ $0.12\pm 0.00$ $0.29\pm 0.04$ $0.07\pm 0.02$
0.200000 96.34 0.99 $0.93\pm 0.04$ $0.04\pm 0.04$ $0.11\pm 0.10$ $0.63\pm 0.28$
0.050000 94.36 0.99 $0.86\pm 0.01$ $0.11\pm 0.00$ $0.24\pm 0.01$ $0.15\pm 0.01$
0.010000 70.88 0.87 $0.95\pm 0.00$ $0.02\pm 0.00$ $0.04\pm 0.00$ $0.84\pm 0.01$
0.005000 52.52 0.84 $0.96\pm 0.01$ $0.02\pm 0.00$ $0.03\pm 0.00$ $0.82\pm 0.01$
0.001000 19.39 0.54 $0.95\pm 0.00$ $0.02\pm 0.00$ $0.04\pm 0.00$ $0.77\pm 0.01$
Table 14: Accuracy–manipulation trade-off for the Gradient Slingshots attack on the “cat” output neuron in VGG9 . Reported are test accuracy (in %), AUROC for the “cat” class, and the mean ± standard deviation of similarity metrics between the manipulated FV and the target image.
$\alpha$ Acc. AUROC CLIP $\uparrow$ MSE $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$
Original 86.33 0.97 $0.74\pm 0.02$ $0.07\pm 0.01$ $0.25\pm 0.02$ $0.03\pm 0.02$
0.9900 86.10 0.97 $0.74\pm 0.02$ $0.08\pm 0.01$ $0.25\pm 0.02$ $0.03\pm 0.02$
0.8000 87.20 0.97 $0.73\pm 0.02$ $0.09\pm 0.01$ $0.25\pm 0.02$ $0.03\pm 0.02$
0.5000 83.90 0.97 $0.78\pm 0.02$ $0.01\pm 0.00$ $0.09\pm 0.01$ $0.20\pm 0.05$
0.1000 86.40 0.97 $0.76\pm 0.02$ $0.01\pm 0.00$ $0.15\pm 0.02$ $0.24\pm 0.04$
0.0500 85.10 0.96 $0.75\pm 0.02$ $0.02\pm 0.00$ $0.11\pm 0.02$ $0.25\pm 0.04$
0.0250 85.15 0.96 $0.79\pm 0.01$ $0.02\pm 0.00$ $0.16\pm 0.02$ $0.33\pm 0.04$
0.0100 81.90 0.95 $0.81\pm 0.02$ $0.01\pm 0.00$ $0.14\pm 0.03$ $0.41\pm 0.05$
0.0050 82.20 0.94 $0.88\pm 0.04$ $0.01\pm 0.00$ $0.08\pm 0.02$ $0.52\pm 0.05$
0.0010 72.70 0.86 $0.87\pm 0.04$ $0.01\pm 0.00$ $0.09\pm 0.02$ $0.54\pm 0.04$
0.0001 9.50 0.54 $0.81\pm 0.02$ $0.02\pm 0.00$ $0.14\pm 0.02$ $0.31\pm 0.07$
Table 15: Accuracy–manipulation trade-off for the Gradient Slingshots attack on the “gondola” output neuron in ResNet-18 . Reported are test accuracy (in %), AUROC for the “gondola” class, and the mean ± standard deviation of similarity metrics between the manipulated FV and the target image.
$\alpha$ Acc. AUROC CLIP $\uparrow$ MSE $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$
Original 71.89 0.99 $0.64\pm 0.03$ $0.09\pm 0.01$ $0.60\pm 0.04$ $0.04\pm 0.02$
0.997 69.50 0.97 $0.73\pm 0.03$ $0.07\pm 0.01$ $0.34\pm 0.04$ $0.06\pm 0.03$
0.995 71.40 0.97 $0.75\pm 0.03$ $0.07\pm 0.01$ $0.34\pm 0.06$ $0.07\pm 0.03$
0.993 68.40 0.96 $0.73\pm 0.03$ $0.07\pm 0.01$ $0.40\pm 0.06$ $0.07\pm 0.03$
0.990 69.30 0.94 $0.73\pm 0.03$ $0.07\pm 0.01$ $0.42\pm 0.05$ $0.07\pm 0.03$
0.950 39.10 0.87 $0.69\pm 0.03$ $0.10\pm 0.01$ $0.58\pm 0.04$ $0.04\pm 0.02$
0.900 11.00 0.79 $0.70\pm 0.02$ $0.08\pm 0.01$ $0.57\pm 0.04$ $0.08\pm 0.03$
0.500 0.60 0.63 $0.62\pm 0.03$ $0.15\pm 0.01$ $0.51\pm 0.03$ $0.03\pm 0.02$
0.100 0.80 0.51 $0.67\pm 0.04$ $0.10\pm 0.04$ $0.61\pm 0.14$ $0.10\pm 0.02$
Table 16: Accuracy–manipulation trade-off for the Gradient Slingshots attack on the “broccoli” output neuron in ViT-L/32 . Reported are test accuracy (in %), AUROC for the “broccoli” class, and the mean ± standard deviation of similarity metrics between the manipulated FV and the target image.
$\alpha$ Acc. AUROC CLIP $\uparrow$ MSE $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$
Original 76.97 1.00 $0.52\pm 0.02$ $0.08\pm 0.00$ $0.74\pm 0.02$ $0.08\pm 0.01$
0.999999 76.58 1.00 $0.51\pm 0.02$ $0.08\pm 0.00$ $0.75\pm 0.03$ $0.07\pm 0.01$
0.999900 76.64 1.00 $0.52\pm 0.03$ $0.09\pm 0.00$ $0.76\pm 0.03$ $0.07\pm 0.01$
0.999500 76.06 0.99 $0.50\pm 0.02$ $0.07\pm 0.00$ $0.77\pm 0.03$ $0.08\pm 0.01$
0.999000 75.96 0.99 $0.49\pm 0.02$ $0.07\pm 0.00$ $0.74\pm 0.02$ $0.09\pm 0.01$
0.990000 72.42 0.99 $0.49\pm 0.02$ $0.06\pm 0.00$ $0.73\pm 0.02$ $0.11\pm 0.01$
0.100000 0.10 0.50 $0.49\pm 0.02$ $0.05\pm 0.00$ $0.76\pm 0.02$ $0.18\pm 0.01$
@@
D.2 Impact of Target Image on Manipulation
@@
We provide additional quantitative evaluation of the effect of the target image on manipulation success, extending the results from Sec. D.2 to an additional set of target images, shown in Fig. 14.
Figure 14: Manipulation results for different target images by source. Shown are example FV outputs, change in overall accuracy (p.p.), and mean±standard deviation of perceptual distance to the target (LPIPS). Manipulation is most effective with natural images.
D.3 Attack Detection
@@
We provide additional qualitative results for our attack detection scheme in Fig. 15, extending the evaluation described in Sec. 6 to additional experimental settings.

### v4

@@ D.5 Accuracy – Manipulation Trade-Off
We expand on the experimental results of the accuracy–manipulation trade-off analysis for the Gradient Slingshots attack, introduced in Sec.˜4.2, with additional findings reported in Tables˜17, 18, 19, 20, 21, 15, 16, 17 and 18
@@
Table 17: Accuracy–manipulation trade-off for the GS attack on the “zero” output neuron in 6-layer CNN . Reported are test accuracy (in %), AUROC for the “zero” class, and the mean ± standard deviation of alignment and similarity metrics, computed over 100 independent FV runs.
Alignment Target Similarity
$\alpha$ AUC Acc. Target Lbl. $\uparrow$ GT Lbl. $\downarrow$ CLIP $\uparrow$ MSE $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$
Original $\mathbf{1.00}$ $\mathbf{99.67}$ $0.29\pm 0.01$ $0.27\pm 0.00$ $0.88\pm 0.01$ $0.13\pm 0.01$ $15\pm 0.01 0.0$$04\pm 0.03$
0.990 $\mathbf{1.00}$ 99.53 $0.29\pm 0.01$ $0.27\pm 0.00$ $0.88\pm 0.01$ $0.13\pm 0.01$ $15\pm 0.01 0.0$$06\pm 0.03$
0.900 $\mathbf{1.00}$ 99.28 $0.30\pm 0.01$ $0.26\pm 0.01$ $0.89\pm 0.01$ $0.10\pm 0.00$ $0.15\pm 0.01$ $0.13\pm 0.02$
0.800 $\mathbf{1.00}$ 98.98 $0.32\pm 0.01$ $0.25\pm 0.00$ $93\pm 0.01 0.0$$03\pm 0.00$ $0.11\pm 0.02$ $0.75\pm 0.01$
[…]
0.500 $\mathbf{1.00}$ 98.10 $0.25\pm 0.01$ $\mathbf{0.24\pm 0.01}$ $0.86\pm 0.01$ $0.12\pm 0.00$ $29\pm 0.04 0.0$$07\pm 0.02$
0.200 0.99 96.34 $0.31\pm 0.03$ $0.25\pm 0.01$ $93\pm 0.04 0.0$$04\pm 0.04$ $0.11\pm 0.10$ $0.62\pm 0.29$
0.050 0.99 94.36 $0.27\pm 0.01$ $\mathbf{0.24\pm 0.00}$ $0.86\pm 0.01$ $0.11\pm 0.00$ $0.24\pm 0.01$ $0.15\pm 0.01$
[…]
0.005 0.84 52.52 $\mathbf{0.33\pm 0.00}$ $0.25\pm 0.00$ $\mathbf{0.96\pm 0.01}$ $\mathbf{0.02\pm 0.00}$ $\mathbf{0.03\pm 0.00}$ $0.82\pm 0.01$
[…]
@@
Table 18: Accuracy–manipulation trade-off for the GS attack on the “cat” output neuron in VGG9 . Reported are test accuracy (in %), AUROC for the “cat” class, and the mean ± standard deviation of alignment and similarity metrics, computed over 100 independent FV runs.
Alignment Target Similarity
$\alpha$ AUC Acc. Target Lbl. $\uparrow$ GT Lbl. $\downarrow$ CLIP $\uparrow$ MSE $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$
Original $\mathbf{0.97}$ $\mathbf{86.33}$ $0.24\pm 0.01$ $0.29\pm 0.01$ $74\pm 0.02 0.0$$08\pm 0.00$ $25\pm 0.02 0.0$$03\pm 0.02$
0.9900 $\mathbf{0.97}$ 86.10 $0.24\pm 0.01$ $0.29\pm 0.01$ $74\pm 0.02 0.0$$08\pm 0.01$ $25\pm 0.02 0.0$$03\pm 0.02$
0.8000 $\mathbf{0.97}$ 87.20 $0.24\pm 0.01$ $0.29\pm 0.01$ $73\pm 0.03 0.0$$09\pm 0.01$ $26\pm 0.02 0.0$$02\pm 0.02$
[…]
0.0500 0.96 85.10 $0.23\pm 0.01$ $0.24\pm 0.01$ $75\pm 0.02 0.0$$02\pm 0.00$ $0.11\pm 0.02$ $0.27\pm 0.05$
0.0250 0.96 85.15 $0.24\pm 0.00$ $0.24\pm 0.01$ $78\pm 0.01 0.0$$02\pm 0.00$ $0.16\pm 0.02$ $0.33\pm 0.04$
[…]
0.0050 0.94 82.20 $\mathbf{0.30\pm 0.04}$ $0.23\pm 0.01$ $\mathbf{0.87\pm 0.04}$ $\mathbf{0.01\pm 0.00}$ $\mathbf{0.08\pm 0.02}$ $0.51\pm 0.05$
[…]
0.0001 0.54 9.50 $0.24\pm 0.01$ $0.24\pm 0.01$ $80\pm 0.01 0.0$$02\pm 0.00$ $0.14\pm 0.02$ $0.30\pm 0.05$
@@
Table 19: Accuracy–manipulation trade-off for the GS attack on the “gondola” output neuron in ResNet-18 . Reported are test accuracy (in %), AUROC for the “gondola” class, and the mean ± standard deviation of alignment and similarity metrics, computed over 100 independent FV runs.
Alignment Target Similarity
$\alpha$ AUC Acc. Target Lbl. $\uparrow$ GT Lbl. $\downarrow$ CLIP $\uparrow$ MSE $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$
Original $\mathbf{0.99}$ $\mathbf{71.89}$ $0.21\pm 0.01$ $0.25\pm 0.02$ $64\pm 0.03 0.0$$09\pm 0.01$ $59\pm 0.04 0.0$$03\pm 0.02$
0.999000 0.98 70.30 $0.21\pm 0.01$ $0.25\pm 0.02$ $65\pm 0.03 0.0$$09\pm 0.01$ $58\pm 0.05 0.0$$04\pm 0.02$
[…]
0.995000 0.97 71.40 $\mathbf{0.25\pm 0.02}$ $0.24\pm 0.02$ $\mathbf{0.75\pm 0.03}$ $\mathbf{0.07\pm 0.01}$ $\mathbf{0.34\pm 0.06}$ $0.07\pm 0.03$
[…]
0.950000 0.87 39.10 $0.22\pm 0.01$ $\mathbf{0.18\pm 0.01}$ $0.69\pm 0.03$ $0.10\pm 0.01$ $58\pm 0.04 0.0$$04\pm 0.02$
0.900000 0.79 11.00 $0.23\pm 0.01$ $0.19\pm 0.01$ $70\pm 0.02 0.0$$08\pm 0.01$ $57\pm 0.03 0.0$$09\pm 0.03$
0.500000 0.63 0.60 $0.23\pm 0.01$ $0.20\pm 0.01$ $0.61\pm 0.03$ $0.15\pm 0.01$ $50\pm 0.03 0.0$$03\pm 0.02$
0.100000 0.51 0.80 $0.22\pm 0.01$ $\mathbf{0.18\pm 0.02}$ $68\pm 0.03 0.0$$09\pm 0.03$ $0.63\pm 0.14$ $\mathbf{0.10\pm 0.02}$
@@
Table 20: Accuracy–manipulation trade-off for the GS attack on the “wolf spider” output neuron in ResNet-50. Extended version of Table˜1 including additional $\alpha$ values and similarity metrics.
Alignment Target Similarity
$\alpha$ AUC Acc. Target Lbl. $\uparrow$ GT Lbl. $\downarrow$ CLIP $\uparrow$ MSE $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$
Original $\mathbf{1.00}$ $\mathbf{76.13}$ $0.23\pm 0.01$ $0.29\pm 0.02$ $0.53\pm 0.02$ $0.12\pm 0.01$ $69\pm 0.01 0.0$$05\pm 0.01$
0.90 $\mathbf{1.00}$ 76.07 $0.22\pm 0.01$ $0.28\pm 0.02$ $0.53\pm 0.02$ $0.12\pm 0.01$ $69\pm 0.01 0.0$$05\pm 0.01$
0.80 $\mathbf{1.00}$ 76.00 $0.22\pm 0.01$ $0.28\pm 0.02$ $0.52\pm 0.02$ $0.13\pm 0.01$ $68\pm 0.01 0.0$$05\pm 0.01$
0.70 $\mathbf{1.00}$ 75.77 $0.23\pm 0.02$ $0.27\pm 0.02$ $0.53\pm 0.03$ $0.12\pm 0.01$ $67\pm 0.02 0.0$$05\pm 0.01$
0.64 $\mathbf{1.00}$ 75.13 $0.31\pm 0.01$ $0.23\pm 0.01$ $0.69\pm 0.02$ $0.12\pm 0.01$ $\mathbf{0.59\pm 0.02}$ $0.04\pm 0.01$
0.62 $\mathbf{1.00}$ 74.83 $\mathbf{0.32\pm 0.01}$ $0.23\pm 0.01$ $0.68\pm 0.02$ $0.12\pm 0.00$ $\mathbf{0.59\pm 0.02}$ $0.04\pm 0.01$
0.60 $\mathbf{1.00}$ 74.51 $\mathbf{0.32\pm 0.01}$ $0.23\pm 0.01$ $0.69\pm 0.02$ $0.12\pm 0.00$ $60\pm 0.02 0.0$$04\pm 0.01$
0.58 $\mathbf{1.00}$ 74.45 $\mathbf{0.32\pm 0.01}$ $0.23\pm 0.01$ $0.69\pm 0.02$ $0.12\pm 0.01$ $60\pm 0.02 0.0$$04\pm 0.01$
0.56 $\mathbf{1.00}$ 73.67 $\mathbf{0.32\pm 0.01}$ $\mathbf{0.22\pm 0.01}$ $0.70\pm 0.02$ $0.11\pm 0.00$ $61\pm 0.01 0.0$$04\pm 0.01$
0.50 $\mathbf{1.00}$ 71.52 $\mathbf{0.32\pm 0.01}$ $0.23\pm 0.01$ $\mathbf{0.72\pm 0.02}$ $0.11\pm 0.00$ $63\pm 0.02 0.0$$05\pm 0.01$
0.40 $\mathbf{1.00}$ 66.58 $0.31\pm 0.01$ $0.24\pm 0.01$ $0.66\pm 0.03$ $0.12\pm 0.01$ $66\pm 0.01 0.0$$05\pm 0.01$
0.30 0.99 52.09 $0.31\pm 0.01$ $0.24\pm 0.01$ $0.65\pm 0.03$ $0.13\pm 0.01$ $65\pm 0.02 0.0$$06\pm 0.01$
[…]
0.10 0.90 30.19 $0.29\pm 0.01$ $0.24\pm 0.01$ $0.59\pm 0.02$ $0.10\pm 0.00$ $71\pm 0.02 0.0$$08\pm 0.02$
[…]
Table 21: Accuracy–manipulation trade-off for the GS attack on the “broccoli” output neuron in ViT-L/32 . Reported are test accuracy (in %), AUROC for the “broccoli” class, and the mean ± standard deviation of alignment and similarity metrics, computed over 100 independent FV runs.
Alignment Target Similarity
$\alpha$ AUC Acc. Target Lbl. $\uparrow$ GT Lbl. $\downarrow$ CLIP $\uparrow$ MSE $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$
Original $\mathbf{1.00}$ $\mathbf{76.97}$ $0.23\pm 0.02$ $0.28\pm 0.02$ $51\pm 0.02 0.0$$08\pm 0.00$ $73\pm 0.03 0.0$$08\pm 0.01$
0.999999 $\mathbf{1.00}$ 76.58 $0.22\pm 0.02$ $0.28\pm 0.02$ $51\pm 0.02 0.0$$08\pm 0.00$ $73\pm 0.03 0.0$$08\pm 0.01$
0.999900 $\mathbf{1.00}$ 76.64 $\mathbf{0.29\pm 0.03}$ $0.27\pm 0.03$ $\mathbf{0.53\pm 0.03}$ $0.09\pm 0.00$ $74\pm 0.04 0.0$$08\pm 0.01$
0.999500 0.99 76.06 $\mathbf{0.29\pm 0.02}$ $\mathbf{0.25\pm 0.01}$ $50\pm 0.02 0.0$$07\pm 0.00$ $74\pm 0.02 0.0$$08\pm 0.01$
0.999000 0.99 75.96 $0.28\pm 0.02$ $\mathbf{0.25\pm 0.01}$ $49\pm 0.02 0.0$$07\pm 0.00$ $72\pm 0.02 0.0$$09\pm 0.01$
0.990000 0.99 72.42 $\mathbf{0.29\pm 0.02}$ $\mathbf{0.25\pm 0.01}$ $49\pm 0.02 0.0$$06\pm 0.00$ $\mathbf{0.71\pm 0.01}$ $0.11\pm 0.01$
[…]
@@
D.6 Impact of Target Image on Manipulation
@@
We provide additional quantitative evaluation of the effect of the target image on manipulation success, extending the results from Sec.˜4.4 to an additional set of target images, shown in Fig.˜19.
Figure 19: Manipulation results for different target images by source. Extended version of Table˜13 including additional target images.
D.7 Attack Detection
@@
We provide additional qualitative results for our attack detection scheme in Fig.˜20, extending the evaluation described in Sec.˜6 to additional experimental settings.